OpenStudy (anonymous):

Converting binary to decimal (step by step) Adding binary to get their sum (step by step) The foundation of number-base How many bits does it take to represent a number, what does it mean?

OpenStudy (unklerhaukus):

\[10011_2=1\times2^4+0\times2^3+0\times2^2+1\times2^1+1\times2^0\\ \qquad\quad=1\times16+0\times8+0\times4+1\times2+1\times1\\ \qquad\quad=16+2+1\\ \qquad\quad=19\]

OpenStudy (mandre):

Sum: 1001 + 1011 = 1001 1011 Working from the right: 1 + 1 = 10 (binary) You get 0 and carry the 1. 0 0+1+1 = 10 You get 0 and carry the 1. 0 0+0+1 = 1 1 1+1 = 10 You get 0 and carry the 1. 0 1 Answer: 10100